# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ 2. Semidiscrete inverse random source problem
$\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,C\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|\mathbf{F}|^{2}dt+\mathbb{E}\int_{Q}\,e^{2s\varphi}|v|^{2}\,dt+\mathcal{O}((sh)^{2})\int_{Q}e^{2s\varphi}|v|^{2}\,dt\right.\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|A_{i}D_{i}v|^{2}+\frac{h}{2}\int_{0}^{T}\int_{\Gamma^{+}}|v|^{2}t_{r}^{i}(e^{2s\varphi})+\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}v|^{2}\right)\,dt\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi})|\eta_{i}|^{2}\,dt+\mathbb{E}\int_{Q}e^{2s\varphi}|v|^{2}\,dt\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+}))}.\end{split}$
@@
Now, by (3.19), we have $\overline{G}_{0}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}}\right)\subset G^{(4)}$ and then
(3.26) $\overline{\mathcal{M}_{0}}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}}\right)\subset\mathcal{M}^{(4)}\subset\mathcal{M}$
for sufficiently small $h\in(0,h_{0})$. Then, observing (3.21) and (3.26) we see that
(3.27) $\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\geq\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D^{2}_{ij}v|^{2}\,dt\\ &\geq C\tau^{-1}e^{2\mu_{4}\tau}\int_{t_{0}-\frac{\varepsilon}{\sqrt{N}}}^{t_{0}+\frac{\varepsilon}{\sqrt{N}}}\left(\mathbb{E}\int_{\mathcal{M}_{0}}|w|^{2}+\sum_{i=1}^{n}\mathbb{E}\int_{(\mathcal{M}_{0})_{i}^{\ast}}\,|D_{i}w|^{2}+\sum_{i,j=1}^{n}\mathbb{E}\int_{(\mathcal{M}_{0})_{ij}^{\ast}}|D^{2}_{i}w|^{2}\right)\,dt.\end{split}$
@@
$Q^{{}^{\prime}}\triangleq\left\{(x,t)\in Q\;:\quad\Phi(x,t)>\mu_{3}+\frac{1}{2}(\mu_{4}-\mu_{3})\right\}.$
@@
(3.29) $\begin{split}&\mathbb{E}\int_{Q}e^{2s\varphi}|\mathbf{F}|^{2}\,dt\\ &\leq Ce^{2\tau\left(\mu_{3}+\frac{1}{2}(\mu_{4}-\mu_{3})\right)}\mathbb{E}\int_{Q\setminus Q^{{}^{\prime}}}\left(|w|^{2}+\sum_{i=1}^{n}\left[\left(1+\frac{h^{4}}{16}\right)|A_{i}D_{i}w|^{2}+\frac{h^{4}}{16}|D_{i}^{2}w|^{2}\right]\right)\,dt\\ &\leq Ce^{\tau(\mu_{4}+\mu_{3})}\|w\|^{2}_{L_{\mathcal{F}}^{2}(0,T;H_{h}^{2}(\mathcal{M}))},\end{split}$
@@
Consider $f(\tau)=ae^{-\frac{1}{2}\tau(\mu_{4}-\mu_{3})}+be^{2\mathcal{C}\tau}$ with $\tau_{1}<\tau<C\varepsilon_{0}h^{-1}$, $a=\|w\|^{2}_{L^{2}_{\mathcal{F}}{(0,T;H_{h}^{2}(\mathcal{M}))}}$ and $b=\|\eta\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma^{+}))}+\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma^{+}))}.$ As $\mu_{4}-\mu_{3}>0$, then from Lemma 3.39 in LOPD the function $f$ satisfies the following.
@@
for any $\tau_{\ast}\leq\frac{2}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\ln{\left(\frac{a}{b}\right)}$. Given that (3.30) is valid for all $\tau\in[\tau_{1},C\varepsilon_{0}h^{-1}]$ and taking $\tau_{\ast}=\frac{2}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\ln{\left(\frac{a}{b}\right)}$, it follows in particular that:
(3.31) $\begin{split}\|w\|^{2}_{L^{2}_{\mathcal{F}}\left((t_{0}-\frac{\varepsilon}{\sqrt{N}},t_{0}+\frac{\varepsilon}{\sqrt{N}};H_{h}^{2}(\mathcal{M}_{0})\right)}\leq\;2\max\left\{be^{2\mathcal{C}\tau_{1}},ae^{-\frac{1}{2h}\mathcal{C}\varepsilon_{0}(\mu_{4}-\mu_{3})},a^{\kappa}b^{1-\kappa}\right\}\end{split}$
@@
$\kappa\triangleq\frac{4\mathcal{C}}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\in(0,1)$
Finally, for the estimate in (3.31), we choose points $t_{0}=\sqrt{2}\varepsilon+\frac{j\varepsilon}{\sqrt{N}}$, for $j=0,1,2,\ldots,m$ where $m$ is chosen such that
$\sqrt{2}\varepsilon+\frac{m\varepsilon}{\sqrt{N}}\leq T-\sqrt{2}\leq\sqrt{2}\varepsilon+\frac{m\varepsilon}{\sqrt{N}}$
Summing the estimates over $j$ from 0 to $m$ and then replacing $\varepsilon$ by $\sqrt{2}\varepsilon$ (which is permissible since $\varepsilon>0$ is arbitrary) produces the desired inequality (1.9), thus completing the proof of Theorem 1.3.
@@ Acknowledgments
R. Lecaros was partially supported by FONDECYT (Chile) Grant 1221892 and Proyecto Interno USM 2025 PI_\_LIR_\_25_\_14. A. A. Pérez acknowledges the support of Vicerrectoría de Investigación y postgrado, Universidad del Bío-Bío, proyect IN2450902 and FONDECYT Grant 11250805. M. F. Prado gratefully acknowledges the support of the Institutional Scholarship Fund of the Universidad de Valparaiso (FIB-UV) and Proyecto Interno USM 2025 PI_\_LIR_\_25_\_14.
Appendix A Calculus results related to the weight functions and Discrete spaces
This section is devoted to presenting key results on discrete operations and discrete spaces, which are essential for the proof of the Carleman estimate. Estimates for the Carleman weight functions in arbitrary dimensions have been established in LOP-2020 . For convenience, we set $r=e^{s\varphi}$ and $\rho=r^{-1}$, where $\varphi(t,x)=\theta(t)\psi(x)$, and consider $s$ and $h$ to be positive parameters with $s$ large and $h$ small. We are particularly interested in the dependence on $s$ and $h$ in the following basic estimates over the domain $G$.
Concerning discrete spaces, we also recall a Sobolev-type inequality from EstimateCarlemanStochasticControl , which will play a key role in controlling norms when switching between discrete spaces.
We assume $s\geq 1$ throughout this section. We begin by presenting two fundamental estimates for the weight function.
Lemma A.1 ((boyer-2010-elliptic, , Lemma 3.7)).
Let $\alpha$ and $\beta$ multi-indices. We have
(A.1) $\begin{split}\partial^{\beta}(r\partial^{\alpha}\rho)=&|\alpha|^{|\beta|}(-s\varphi)^{|\alpha|}\lambda^{|\alpha+\beta|}(\nabla\psi)^{\alpha+\beta}+|\alpha||\beta|(s\varphi)^{|\alpha|}\lambda^{|\alpha+\beta|-1}\mathcal{O}(1)+s^{|\alpha|-1}|\alpha|(|\alpha|-1)\mathcal{O}_{\lambda}(1)\\ =&\mathcal{O}_{\lambda}(s^{|\alpha|}).\end{split}$
Let $\sigma\in[-1,1]$ and $i\in\llbracket 1,d\rrbracket$. We have
(A.2) $\begin{split}\partial^{\beta}\left(r(x)(\partial^{\alpha}\rho)(x+\sigma h_{i})\right)=\mathcal{O}_{\lambda,\varepsilon}(s^{|\alpha|}(1+(sh)^{|\beta|}))e^{\mathcal{O}_{\lambda}(sh)}.\end{split}$
Provided $sh\leq\varepsilon$ we have $\partial^{\beta}\left(r(x)(\partial^{\alpha}\rho)(x+\sigma h_{i})\right)=\mathcal{O}_{\lambda,\varepsilon}(s^{|\alpha|})$. The same expressions hold with $r$ and $\rho$ interchanged and with $s$ changed to s-s.
Another useful estimate is the following
Corollary A.2 ((boyer-2010-elliptic, , Corollary 3.8)).
Let $\alpha,\beta$ and $\delta$ multi-indices. We have
(A.3) $\begin{split}\partial^{\delta}\left(r^{2}\,\partial^{\alpha}\rho\,\partial^{\beta}\rho\right)=&|\alpha+\beta|^{|\delta|}(-s\varphi)^{|\alpha+\beta|}\lambda^{|\alpha+\beta+\delta|}(\nabla\psi)^{\alpha+\beta+\delta}\\ &+|\delta||\alpha+\beta|(s\varphi)^{|\alpha+\beta|}\lambda^{|\alpha+\beta+\delta|-1}\mathcal{O}(1)\\ &+s^{|\alpha+\beta|-1}\left(|\alpha|(|\alpha|-1)+|\beta|(|\beta|-1)\right)\mathcal{O}(1)\\ &=\mathcal{O}_{\lambda}(s^{|\alpha+\beta|}).\end{split}$
Mimicking the notation for multi-index in the continuous setting, we define for the 2-dimensional index $k=(k_{i},k_{j})$ with $k_{i},k_{j}\in\mathbb{N}$ the discrete operator $\mathbf{D}_{h}^{k}:=D_{i}^{k_{i}}D_{j}^{k_{j}}$. Analogously, for the average operator, we define $\mathbf{A}_{h}^{l}:=A_{i}^{l_{i}}A_{j}^{l_{j}}$ with $l:=(l_{i},l_{j})$ being a two-dimensional index.
Proposition A.3 ((AA:perez:2024, , Proposition 3.3)).
Let $\alpha$ be a multi-index and $k,l$ be two-dimensional multi-indices, respectively. Let $i,j\in\llbracket 1,d\rrbracket$; provided that $sh\leq\varepsilon$, we have
(A.4) $r\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\alpha}\rho$ $=r\partial^{k+\alpha}\rho+s^{|\alpha|+k_{i}}\mathcal{O}_{\lambda}((sh)^{2})$
$\quad+s^{|\alpha|+k_{j}}\mathcal{O}_{\lambda}((sh)^{2})+s^{|\alpha+k|}\mathcal{O}_{\lambda}((sh)^{2}).$
Another useful estimate is the following lemma
Lemma A.4 ((AA:perez:2024, , Lemma 3.4)).
Let $\alpha$, $\beta$ and $\delta$ be multi-indices and let $k,l$ be two-dimensional multi-indices. Provided $sh\leq 1$, we have
(A.5) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\left(\partial^{\beta}\left(r\partial^{\alpha}\rho\right)\right)=\partial_{i}^{k_{i}}\partial_{j}^{k_{j}}\partial^{\beta}\left(r\partial^{\alpha}\rho\right)+h^{2}\mathcal{O}_{\lambda}(s^{|\alpha|}).$
Moreover, for $i\neq j\in\llbracket 1,d\rrbracket$ we have
(A.6) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\delta}\left(r^{2}\left(\partial^{\alpha}\rho\right)\partial^{\beta}\rho\right)=\partial^{k+\delta}\left(r^{2}(\partial^{\alpha}\rho)\partial^{\beta}\rho\right)+h^{2}\mathcal{O}(s^{|\alpha|+|\beta|}),$
and for $\sigma,\sigma^{\prime}\in[0,1]$ we get
(A.7) $\mathbf{A}_{h}^{l}\mathbf{D}^{k}_{h}\partial^{\beta}\left(r(x)\partial^{\alpha}\rho(x+\sigma he_{i}\right)$ $=\mathcal{O}_{\lambda}(s^{|\alpha|}),$
(A.8) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\delta}\left(r^{2}(x)\left(\partial^{\alpha}\rho(x+\sigma he_{i})\right)\partial^{\beta}\rho(x+\sigma^{\prime}he_{j})\right)$ $=\mathcal{O}_{\lambda}(s^{|\alpha|+|\beta|}).$
Moreover, we present estimates for various compositions of the discrete average and derivative operators applied to the Carleman weight function. These results will play a fundamental role in the proof of the semidiscrete Carleman estimate.
Theorem A.5 ((AA:perez:2024, , Theorem 3.5)).
Let $\alpha$ be a multi-index and $l,k,n,m,p$ and $q$ two-dimensional multi-indices. For $i,j\in\llbracket 1,d\rrbracket$, and for $sh\leq 1$, we have
(A.9) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\alpha}(r\mathbf{A}_{h}^{m}\mathbf{D}_{h}^{n}\rho)=$ $\partial^{n+\alpha}(r\partial^{n}\rho)++s^{|\alpha|+n_{i}}\mathcal{O}_{\lambda}((sh)^{2})$
$+s^{|\alpha|+n_{j}}\mathcal{O}_{\lambda}((sh)^{2})+s^{|n|}\mathcal{O}_{\lambda}((sh)^{2})$
and
(A.10) $\mathbf{A}_{h}^{l}\mathbf{D}_{h}^{k}\partial^{\beta}(r^{2}\mathbf{A}^{p}_{h}\mathbf{D}^{q}_{h}(\partial^{\alpha}\rho)\mathbf{A}_{h}^{m}\mathbf{D}_{h}^{n}\rho)=$ $\partial^{k+\beta}(r^{2}(\partial^{q+\alpha}\rho)\partial^{n}\rho)$
$+s^{|\alpha+n+q|}\mathcal{O}_{\lambda}((sh)^{2}).$
Finally, we conclude this section with a discrete Sobolev-type inequality, which plays a key role in controlling norms when transferring between discrete spaces.
Theorem A.6 ((EstimateCarlemanStochasticControl, , Theorem 1.5)).
Let $n>1$. There exists a constant $C$ independent of $h$ such that the following inequality holds:
$\|u\|_{L_{h}^{p^{\ast}}(\mathcal{M)}}\leq C\|u\|_{W_{h}^{1,p}(\mathcal{M})}\quad\forall u\in W_{h}^{1,p}(\mathcal{M}),$
with $u=0$ on $\partial\mathcal{M}$, where
  • If $1\leq p<n$, then $p^{\ast}$ is given by $\frac{1}{p^{\ast}}=\frac{1}{p}-\frac{1}{n}$.
  • If $p=n$ then $p^{\ast}\in[p,\infty)$.
@@ Appendix A Calculus results related to the weight functions and Discrete spaces
Appendix B Proof of Semidiscrete Carleman estimate with homogeneous Dirichlet condition (Theorem 1.6).
@@
In this section, we prove the semidiscrete Carleman estimate under homogeneous Dirichlet conditions. The argument follows a classical approach (see fursikov-1996 ), based on conjugating the original operator with a suitably chosen exponential weight. Our estimate is closely related to the one presented in boyer-2010-1d-elliptic , and we adopt its methodology whenever possible. The main difference is that we do not impose any boundary condition on the state or on the weight function, which leads to additional boundary terms that will be handled later on.
@@
The proof is divided into three steps. First, in Section B.1, we decompose the conjugated operator into its symmetric part $P_{1}$, antisymmetric part $P_{2}$, and an auxiliary term $R$, which helps simplify the calculations. Next, in Section B.2, we estimate the cross-inner products between these operators. Finally, in Section B.3, we return to the original variable to conclude the argument.
B.1. Conjugated operator

### v2

@@ 2. Semi-discrete inverse random source problem
$\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\leq\,\mathcal{C}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|\mathbf{F}|^{2}dt+\mathbb{E}\int_{Q}\,e^{2s\varphi}|v|^{2}\,dt+\mathcal{O}((sh)^{2})\int_{Q}e^{2s\varphi}|v|^{2}\,dt\right.\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|A_{i}D_{i}v|^{2}+\frac{h}{2}\int_{0}^{T}\int_{\Gamma}|v|^{2}t_{r}^{i}(e^{2s\varphi})+\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}v|^{2}\right)\,dt\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi})|t_{r}^{i}(D_{i}v)|^{2}\,dt+\mathbb{E}\int_{Q}e^{2s\varphi}|v|^{2}\,dt\right)+\mathcal{C}\tau^{3}e^{\tau\mathcal{C}}\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma))}.\end{split}$
@@
Now, by (3.21), we have $\overline{\mathcal{M}}_{0}\times\left(t_{0}-\frac{\varepsilon}{\sqrt{\tilde{N}}},t_{0}+\frac{\varepsilon}{\sqrt{\tilde{N}}}\right)\subset G^{(4)}\cap Q$ and then, observing (3.23), we see that
(3.28) $\begin{split}&J(v)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}v|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|a_{3}v|^{2}dt\\ &\geq\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|v|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}v|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D^{2}_{ij}v|^{2}\,dt\\ &\geq\mathcal{C}\tau^{-1}e^{2\mu_{4}\tau}\int_{t_{0}-\frac{\varepsilon}{\sqrt{\tilde{N}}}}^{t_{0}+\frac{\varepsilon}{\sqrt{\tilde{N}}}}\left(\mathbb{E}\int_{\mathcal{M}_{0}}|w|^{2}+\sum_{i=1}^{n}\mathbb{E}\int_{(\mathcal{M}_{0})_{i}^{\ast}}\,|D_{i}w|^{2}+\sum_{i,j=1}^{n}\mathbb{E}\int_{(\mathcal{M}_{0})_{ij}^{\ast}}|D^{2}_{i}w|^{2}\right)\,dt.\end{split}$
@@
$Q^{{}^{\prime}}:=\left\{(x,t)\in Q\;:\quad\varphi(x)\theta(t)>\mu_{3}+\frac{1}{2}(\mu_{4}-\mu_{3})\right\}.$
@@
(3.30) $\begin{split}&\mathbb{E}\int_{Q}e^{2s\varphi}|\mathbf{F}|^{2}\,dt\\ &\leq\mathcal{C}e^{2\tau\left(\mu_{3}+\frac{1}{2}(\mu_{4}-\mu_{3})\right)}\mathbb{E}\int_{Q\setminus Q^{{}^{\prime}}}\left(|w|^{2}+\sum_{i=1}^{n}\left[\left(1+\frac{h^{4}}{16}\right)|A_{i}D_{i}w|^{2}+\frac{h^{4}}{16}|D_{i}^{2}w|^{2}\right]\right)\,dt\\ &\leq\mathcal{C}e^{\tau(\mu_{4}+\mu_{3})}\|w\|^{2}_{L_{\mathcal{F}}^{2}(0,T;H_{h}^{2}(\mathcal{M}))},\end{split}$
@@
Consider $f(\tau)=ae^{-\frac{1}{2}\tau(\mu_{4}-\mu_{3})}+be^{2\mathcal{C}\tau}$ with $\tau_{1}<\tau<\mathcal{C}\varepsilon_{0}h^{-1}$, $a=\|w\|^{2}_{L^{2}_{\mathcal{F}}{(0,T;H_{h}^{2}(\mathcal{M}))}}$ and $b=\|{\partial_{\nu}w}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\Gamma))}+\|\xi\|^{2}_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1}(\Gamma))}.$ As $\mu_{4}-\mu_{3}>0$, then from Lemma 3.39 in Lecaros et al. (2023) the function $f$ satisfies the following.
@@
for any $\tau_{\ast}\leq\frac{2}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\ln{\left(\frac{a}{b}\right)}$. Given that (3.31) is valid for all $\tau\in[\tau_{1},\mathcal{C}\varepsilon_{0}h^{-1}]$ and taking $\tau_{\ast}=\frac{2}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\ln{\left(\frac{a}{b}\right)}$, it follows in particular that:
(3.32) $\begin{split}\|w\|^{2}_{L^{2}_{\mathcal{F}}\left((t_{0}-\frac{\varepsilon}{\sqrt{\tilde{N}}},t_{0}+\frac{\varepsilon}{\sqrt{\tilde{N}}};H_{h}^{2}(\mathcal{M}_{0})\right)}\leq\;2\max\left\{be^{2\mathcal{C}\tau_{1}},ae^{-\frac{1}{2h}\mathcal{C}\varepsilon_{0}(\mu_{4}-\mu_{3})},a^{\kappa}b^{1-\kappa}\right\}\end{split}$
@@
$\kappa:=\frac{4\mathcal{C}}{4\mathcal{C}+(\mu_{4}-\mu_{3})}\in(0,1)$
By the argument, we see that the constant of the right-hand side of (3.32) is also independent of $t_{0}$.
Since $t_{0}\in[\sqrt{2}\varepsilon,\,T-\sqrt{2}\varepsilon]$, in (3.32) we choose the points $t_{0}^{\,j}=\sqrt{2}\varepsilon+\frac{j\varepsilon}{\sqrt{\tilde{N}}}$ for $j=0,1,2,\ldots,m$, where $m$ is chosen such that
$\sqrt{2}\varepsilon+\frac{m\varepsilon}{\sqrt{\tilde{N}}}\leq T-\sqrt{2}\varepsilon\leq\sqrt{2}\varepsilon+\frac{(m+1)\varepsilon}{\sqrt{\tilde{N}}}\leq T.$
Summing the corresponding estimates with respect to $j$ from 0 to $m$, and then replacing $\varepsilon$ by $\sqrt{2}\varepsilon$ (which is admissible since $\varepsilon>0$ is arbitrary), we obtain the desired inequality (1.9), and this completes the proof of Theorem 1.3.
4. Comments and concluding remarks
In Theorem 1.1, we establish a Lipschitz stability result for an inverse source problem for semi-discrete stochastic parabolic equations in arbitrary spatial dimensions. The stability is obtained from interior observations together with the terminal value of the solution. In contrast with [20], the observation region is an arbitrary open subset of the spatial domain. Moreover, the regularity assumption on the coefficient $a_{2}$ is weaker in comparison with the existence literature. For this reasons, this problem is new. In Theorem 1.3, we prove a Hölder stability result for a Cauchy problem using the terminal value of the solution and the trace of the spatial derivatives on an arbitrary open subset of the boundary over a time interval. The novelty here lies in the treatment of arbitrary spatial dimensions and in allowing the boundary observation set to be a arbitrary open subset. From a methodological point of view, we derive three new Carleman estimates adapted to the different settings considered in the paper (see Theorems 1.5, 3.3 and 3.7). These estimates require new weight functions and refined analytical arguments in the discrete stochastic framework. In particular, they allow us to establish observability inequalities on arbitrary open subsets, which is not addressed in [20]. See, for instance, assumption 3.1 and the related discussion. Moreover, Theorem 3.3 and 3.7 generalize the results presented in [20] since the measurement on the boundary are arbitrary open set sampled on the mesh.
Several variants of the inverse problem are related to the results presented in this manuscript. In particular, an interesting open problem is to obtain a stability result similar to Theorem 1.1, but for a drift term $f$ in system 1.4, or, in a more general case, for both the drift and diffusion terms at the same time. In this direction, we can refer to the uniqueness results in Lü and Zhang (2024); Lü (2012) obtained in a continuous setting, where one of the main tools is the use of Carleman estimates. Although a similar analysis could be considered in our case, the arguments cannot be applied directly to the semi-discrete framework. In fact, besides Carleman estimates, some additional tools are needed in the continuous case, and it is not clear how to adapt them to the semi-discrete setting. This point is important because uniqueness results for inverse problems are not always true in the semi-discrete context. For example, Theorem 1.1 implies a uniqueness result, while Remark 1.4 shows that uniqueness cannot be obtained from the corresponding stability result. This difference shows that uniqueness is a delicate issue in the semi-discrete case and makes the study of the results in Lü and Zhang (2024); Lü (2012) under this framework interesting.
Furthermore, the result presented in this work could also be applied to extend the recent result provided in Wu et al. (2025). Indeed, in Wu et al. (2025) the authors established Lipschitz stability for an inverse problem of a fully-discrete stochastic hyperbolic equation in 1-D spatial dimension; and their results are based in a new Carleman estimate.
@@ Acknowledgments
R. Lecaros was partially supported by FONDECYT (Chile) Grant 1221892 and Proyecto Interno USM 2025 PI_\_LIR_\_25_\_14. A. A. Pérez acknowledges the support of Vicerrectoría de Investigación y postgrado, Universidad del Bío-Bío, project IN2450902 and FONDECYT Grant 11250805. M. F. Prado gratefully acknowledges the support of the Institutional Scholarship Fund of the Universidad de Valparaiso (FIB-UV) and Proyecto Interno USM 2025 PI_\_LIR_\_25_\_14.
@@
Appendix A Proof of semi-discrete Carleman estimate with homogeneous Dirichlet condition (Theorem 1.5).
@@
In this section, we prove the semi-discrete Carleman estimate under homogeneous Dirichlet conditions. The argument follows a classical approach (see Fursikov and Imanuvilov (1996)), based on conjugating the original operator with a suitably chosen exponential weight. Our estimate is closely related to the one presented in Boyer et al. (2010a), and we adopt its methodology whenever possible. The main difference is that we do not impose any boundary condition on the state or on the weight function, which leads to additional boundary terms that will be handled later on.
@@
The proof is divided into three steps. First, in Section A.1, we decompose the conjugated operator into its symmetric part $P_{1}$, antisymmetric part $P_{2}$, and an auxiliary term $R$, which helps simplify the calculations. Next, in Section A.2, we estimate the cross-inner products between these operators. Finally, in Section A.3, we return to the original variable to conclude the argument.
A.1. Conjugated operator